OpenStudy (anonymous):

HELP WITH THREE QUESTIONS MEDAL AND FAN WILL BE GIVEN A $4,000.00 principal earns 5% interest, compounded annually. After 4 years, what is the balance in the account? (1 point) $500,000.00 $4,862.03 $4,600.00 $20,250.00 24. A $6,000.00 principal earns 8% interest, compounded semiannually. After 35 years, what is the balance in the account? (1 point) $22,800.00 $39,600.00 $88,712.07 $93,429.71 25. A tractor costs $15,450.00 and depreciates in value by 14% per year. How much will the tractor be worth after 3 years? (1 point) $22,8

OpenStudy (anonymous):

@KirbyLegs can you help please

OpenStudy (anonymous):

@oneinmillon do you know

OpenStudy (anonymous):

the first one is the third choice #24 is the second choice #25 i dunno

OpenStudy (anonymous):

$22,889.85 $15,408.00 $8,451.28 $9,827.07 those are the answer choices for 25

OpenStudy (anonymous):

24.  Where P is the present value, or principal r is the compounded rate of interest m is the number of divisions per year (in this case 2, since it's compounded semi-annually, or per half-year) n is the number of years. n is still 35 years okay? And don't forget that r should not be in percent, but in decimal: 8% = 0.08 P is 6000, so... we basically have 6000(1.04)^70 we get 1.04 from 1+r/m r = 0.08 since the rate is 8% m is 2, since it's compounded per half-year (semi-annually) so.. 1+0.082/2=1+0.04=1.04 so, answers D fro question 24